OpenStudy (anonymous):

2/x+4+6/x+4=-4 The solution set is?

OpenStudy (anonymous):

\[x \in \left\{ -2/3 \right\}\]

OpenStudy (anonymous):

multiply thru by x+4 2 + 6 = -4x - 16 - 4x = 24 x = -6 i am assuming that x+4 is the denominator in fractions

OpenStudy (anonymous):

x∈{−6}